# On the Classification of Dillon’s APN Hexanomials

### v1

@@ 7.1. Main results
Our computational searches found 390 and 28,170 APN hexanomials for $\mathbb{F}_{2^{2}}$ and $\mathbb{F}_{2^{4}}$ respectively, classified into at least 10 and 64 distinct CCZ-invariant classes. For $q=2$, 80% of classes are not BC-equivalent to the known Budaghyan-Carlet family; for $q=4$, this rises to 93.75%. This demonstrates that Dillon’s family is significantly richer than previously recognized, validating his 2006 intuition. Notably, none of the discovered APN functions are permutations, suggesting structural incompatibility between this polynomial form and bijectivity.
@@ 7.2. Open questions and future directions
Looking forward, completing the classification for $q\in\{32,64,128\}$ would definitively identify all APN hexanomials in these fields and verify whether our theoretical obstructions extend to all $q\geq 2$. The 9,120 APN functions found for $q=4$ likely contain multiple infinite families; identifying patterns in their coefficients could lead to new constructions generalizing Budaghyan-Carlet.
@@ Acknowledgments
The third-named author (PS) would like to thank the first-named author (DB) for the invitation at the Dipartimento di Matematica e Informatica at Università degli Studi di Perugia, Italy, and the great working conditions while this paper was being started. The first-named author (DB) and second-named author (GGG) thank the Italian National Group for Algebraic and Geometric Structures and their Applications (GNSAGA—INdAM) which supported the research.

### v3

@@ 7.1. Main results
Our computational searches found 390 and 28,170 APN hexanomials for $\mathbb{F}_{2^{2}}$ and $\mathbb{F}_{2^{4}}$ respectively, all belonging to a single CCZ-equivalence class—the known Budaghyan-Carlet family. For larger fields, we found 218 APN functions for $q=8$ in 3 distinct classes and 34 functions for $q=16$ in 2 distinct classes, all outside the BC family. This demonstrates that Dillon’s family is significantly richer than previously recognized, validating his 2006 intuition. Notably, none of the discovered APN functions are permutations, suggesting structural incompatibility between this polynomial form and bijectivity.
@@ 7.2. Open questions and future directions
Looking forward, completing the classification for $q\in\{32,64,128\}$ would definitively identify all APN hexanomials in these fields and verify whether our theoretical obstructions extend to all $q\geq 2$. The discovery of new CCZ-equivalence classes for $q\geq 8$ outside the Budaghyan-Carlet family suggests that larger fields may harbor additional inequivalent constructions.
@@ Acknowledgments
The third-named author (PS) would like to thank the first-named author (DB) for the invitation at the Dipartimento di Matematica e Informatica at Università degli Studi di Perugia, Italy, and the great working conditions while this paper was being started. The first-named author (DB) and second-named author (GGG) thank the Italian National Group for Algebraic and Geometric Structures and their Applications which supported the research (INdAM – GNSAGA Project, CUP E53C24001950001).